OpenStudy (anonymous):

What is the simplified form of y^2 / x all over y^5 / x^2 ?

OpenStudy (anonymous):

\[\frac{ a/b }{ c/d }=\frac{ a }{ b }*\frac{ d }{ c }\] When a fraction is divided by a fraction the problem can be rewritten as the top fraction multiplied by the inverse of the bottom fraction

OpenStudy (anonymous):

I got xy^3 is that wrong?

OpenStudy (anonymous):

i got \[\frac{ x }{ y^3 }\]

OpenStudy (anonymous):

Oh wait you're right. I just did it again. I see what I did wrong. :)

OpenStudy (anonymous):

great!